### v1

# Improving Numerical Stability of Normalized Mutual Information Estimator on High Dimensions

Abstract
Mutual information provides a powerful, general-purpose metric for quantifying the amount of shared information between variables. Estimating normalized mutual information using a k-Nearest Neighbor (k-NN) based approach involves the calculation of the scaling-invariant k-NN radius. Calculation of the radius suffers from numerical overflow when the joint dimensionality of the data becomes high, typically in the range of several hundred dimensions. To address this issue, we propose a logarithmic transformation technique that improves the numerical stability of the radius calculation in high-dimensional spaces. By applying the proposed transformation during the calculation of the radius, numerical overflow is avoided, and precision is maintained. Proposed transformation is validated through both theoretical analysis and empirical evaluation, demonstrating its ability to stabilize the calculation without compromizing the precision of the results.
Index Terms:
Mutual Information, Nonparametric Statistics, Nearest Neighbor Methods, Numerical Stability
I Introduction
Mutual Information (MI) is a fundamental concept in information theory [1, 2, 3], widely applied in fields such as machine learning [4, 5, 6, 7], signal processing [8, 9], and statistics [10, 11, 12]. Unlike traditional metrics like Pearson correlation, MI captures both linear and non-linear dependencies between random variables, providing a more comprehensive measure of association. Accurate MI estimation is therefore critical for a wide range of scientific and engineering applications.
MI quantifies the reduction in uncertainty about one variable given knowledge of another. Formally, for two random variables $X$ and $Y$ with joint distribution $p(x,y)$ and marginals $p(x)$ and $p(y)$, MI is defined as the Kullback-Leibler divergence between the joint distribution and the product of the marginals [3]:
$\phantom{.}I(X;Y)\triangleq D_{KL}\left(p(x,y)\mid\mid p(x)p(y)\right).$ (1)
For discrete random variables, MI is calculated as [1]:
$\phantom{.}I(X;Y)=\sum_{x\in X}\sum_{y\in Y}p(x,y)\ln\frac{p(x,y)}{p(x)p(y)}.$ (2)
Alternatively, MI can be expressed through Shannon entropy [13]:
$\phantom{.}I(X;Y)=H(X)+H(Y)-H(X;Y).$ (3)
For continuous variables, differential entropy is used [13], replacing summations with integrals:
$\phantom{.}I(X;Y)=\int_{x\in X}\!\int_{y\in Y}p(x,y)\ln\frac{p(x,y)}{p(x)p(y)}dx\,dy.$ (4)
Estimating MI for continuous variables poses significant challenges due to the difficulty of accurately estimating joint probability distributions—especially in high-dimensional spaces where the curse of dimensionality [14] leads to data sparsity. Traditional MI estimation methods, such as histogram-based [15, 16] and kernel density estimation [17], tend to perform poorly in higher dimensions. The k-Nearest Neighbors (k-NN) approach [18, 19, 20] is more flexible and widely used, but recent advances in neural-based estimators [21, 22] and variational methods [23, 24] show potential for tackling complex, high-dimensional problems, despite challenges with sparse data and long-tailed distributions [25].
Figure 1: Comparison of the original, proposed, and dominant-term equations of the scaling-invariant k-NN radius in Equation (8), showing the asymptotic behavior of denominator values $V$ as a function of joint dimensionality $d_{X}+d_{Y}$ for fixed radii $\varepsilon$.
In this work, we utilize the Kraskov-Stögbauer-Grassberger (KSG) estimator [19], which estimates MI using (3) and entropy values derived from the k-NN algorithm [26]. To normalize MI and scale it between 0 and 1, we apply the method proposed by Nagel et al. [20], which is invariant under variable transformations and provides an upper bound for MI. However, this normalization method suffers from numerical instability in high-dimensional spaces, especially when computing the scaling-invariant k-NN radius.
To address this, we propose a logarithmic transformation technique that stabilizes the calculation of the scaling-invariant k-NN radius, without sacrificing precision, as illustrated in Figure 1. This proposed enhancement improves the robustness of the estimator for large-scale, high-dimensional datasets, further strengthening its applicability in domains such as computational neuroscience [27], bioinformatics [28], and multivariate statistics [29]—as well as other fields discussed earlier, including machine learning, signal processing, and broader statistical applications.
II Background
The KSG estimator [19] use k-Nearest Neighbors (k-NN) to estimate joint and marginal entropies directly from the data. For each data point, the algorithm calculates the distance to its k-th nearest neighbor in joint space and counts the number of points within that same distance in the marginal spaces. These distances and counts are used to estimate the densities and entropies of the joint and marginal distributions, making the estimator non-parametric and independent of explicit probability distribution assumptions. The KSG estimator is expressed as:
$\begin{split}\widehat{I(X;Y)}=\psi(N)+\psi(k)-\langle\psi(n_{X}+1)+\psi(n_{Y}+1)\rangle,\end{split}$ (5)
[…]
II-A Normalizing Mutual Information
Normalized Mutual Information (NMI), as defined in [20], is the mutual information $I(X;Y)$ normalized by the geometric mean of the marginal entropies $H(X)$ and $H(Y)$:
$\phantom{.}I_{\text{NMI}}(X;Y)\triangleq\frac{I(X;Y)}{\sqrt{H(X)H(Y)}}.$ (6)
However, this formulation applies only to discrete entropies. For continuous variables, differential entropy is not invariant under coordinate transformations, which can lead to negative entropy values. Jaynes [31] introduced relative entropy to resolve these issues, incorporating an invariant measure $m(x)$ to correct for scale and parametrization changes:
$\phantom{.}H_{r}(X)=-\int_{x\in X}p(x)\ln\frac{p(x)}{m(x)}dx.$ (7)
Nagel et al. [20] applied the concept of relative entropy to ensure that MI and its normalization factor in (6) remain consistent across different coordinate systems by using invariant measures derived from local volumes around data points.
II-B Scaling-Invariant k-NN Radius and Entropy Estimation
A critical aspect of Nagel et al.’s method [20] is the calculation of the k-NN radii, $\varepsilon$, which estimates the local densities of data points. To avoid bias from scale or dimensionality, the raw k-NN distances are normalized:
$\phantom{,}\tilde{\varepsilon}=\frac{\varepsilon}{\langle\varepsilon^{d_{X}+d_{Y}}\rangle^{1/(d_{X}+d_{Y})}}=\frac{\varepsilon}{V},$ (8)
where $V=\langle\varepsilon^{d_{X}+d_{Y}}\rangle^{1/(d_{X}+d_{Y})}$ is the normalization factor, and $d_{X}$ and $d_{Y}$ are the dimensionalities of the marginal spaces.
II-C Entropy Estimators with Invariant Measures
Nagel et al. [20] developed entropy estimators that incorporate the scaling-invariant k-NN radii from Equation (8). These estimators use relative entropy to ensure that marginal and joint entropy estimates are invariant across different scales and transformations. The marginal and joint relative entropy estimators are given as:
$\phantom{,}\hat{H}_{r}(X)=-\langle\psi(n_{x}+1)\rangle+\psi(N)+d_{X}\langle\ln\tilde{\varepsilon}\rangle,$ (9)
$\phantom{.}\hat{H}_{r}(X,Y)=-\psi(k)+\psi(N)+(d_{X}+d_{Y})\langle\ln\tilde{\varepsilon}\rangle.$ (10)
NMI is then calculated by substituting these marginal and joint entropy estimates into Equations (3) and (6).
III Proposed transformation
We begin by taking the logarithm of the normalization factor—denominator in (8)—and applying the product and power rules of logarithms, as described in [32, p. 37]:
$\phantom{.}\ln V=\frac{1}{d_{X}+d_{Y}}\left[\ln\dfrac{1}{N}+\ln\sum_{i=1}^{N}{\varepsilon_{i}^{d_{X}+d_{Y}}}\right].$ (11)
Next, the largest k-NN radius $\varepsilon_{\text{max}}$ is identified:
$\phantom{.}\varepsilon_{\text{max}}=\max_{1\leq i\leq N}{\varepsilon_{i}}.$ (12)
Using the log-sum-exp formula from [33, p. 844] and applying the quotient rule of logarithms from [32, p. 37] equation (11) is refactored by factoring out the maximum value $\varepsilon_{\text{max}}$ from (12):
$\phantom{.}\ln V=\ln\varepsilon_{\text{max}}+\tfrac{1}{d_{X}+d_{Y}}\ln\left(\frac{\sum_{i=1}^{N}{\left(\frac{\varepsilon_{i}}{\varepsilon_{\text{max}}}\right)^{d_{X}+d_{Y}}}}{N}\right).$ (13)
Substituting the result from (13) back into the original equation (8)—see Equation (14)—avoids numerical overflow during the calculation of the scaling-invariant radii, as all the terms $\nicefrac{{\varepsilon_{i}}}{{\varepsilon_{\text{max}}}}$ are guaranteed to be less than or equal to one.
$\tilde{\varepsilon}=\varepsilon/{\rm e}^{\ln V}$ (14)
Additionally, for very large values of $d_{X}+d_{Y}$, the scaling-invariant radii calculation is dominated by $\varepsilon_{\text{max}}$—as illustrated in Figure 1—since the second term in Equation (13) becomes negligible. This happens because $\nicefrac{{\varepsilon_{i}}}{{\varepsilon_{\text{max}}}}$ for $\varepsilon_{i}<\varepsilon_{\text{max}}$ decay exponentially to zero as $d_{X}+d_{Y}\to\infty$, effectively eliminating their contribution to the sum. However, the terms where $\varepsilon_{i}=\varepsilon_{\text{max}}$ remain equal to $1$ and continue to contribute to the sum, ensuring it stays finite and preventing the logarithm from tending to minus infinity. Moreover, the factor $\frac{1}{d_{X}+d_{Y}}\to 0$ as $d_{X}+d_{Y}\to\infty$, which further diminishes the impact of the second term. As a result, Equation (14) simplifies to $\tilde{\varepsilon}=\varepsilon/\varepsilon_{\text{max}}$ for large $d_{X}+d_{Y}$.
IV Numerical Experiments
We evaluate the proposed transformation to stabilize calculations while maintaining precision. We compare both the baseline estimator—detailed in Equations (9) and (10)—and our modified version, which incorporates the scaling-invariant radius calculation from Equation (14), across different correlation coefficients and dimensionalities. The results are also benchmarked against the theoretical ground-truth NMI values provided in Equation (20).
IV-A Experimental Setup
We generate pairs $(\mathbf{x}_{i},\mathbf{y}_{i})$ from a correlated Gaussian distribution, where $\mathbf{x}_{i},\mathbf{y}_{i}\in\mathbb{R}^{d}$ represent the i-th sample pair. Here, $\mathbf{x}_{i}$ and $\mathbf{y}_{i}$ are individual samples, while $X$ and $Y$ denote the random variables corresponding to the entire sets of samples.
The correlation between corresponding components of $\mathbf{x}_{i}$ and $\mathbf{y}_{i}$ is defined by the coefficient $\rho$, following the model described by Belghazi et al. [21]:
$\phantom{,}\text{corr}(X^{i},Y^{j})=\delta_{ij}\rho,$ (15)
where $\delta_{ij}$ is the Kronecker delta, ensuring that only corresponding components (i.e., $i=j$) are correlated. The joint covariance matrix for $X$ and $Y$ is:
$\phantom{,}\Sigma=\begin{pmatrix}I_{d}&\rho I_{d}\\ \rho I_{d}&I_{d}\end{pmatrix},$ (16)
where $I_{d}$ is the identity matrix of size $d$. Both $\mathbf{x}_{i}$ and $\mathbf{y}_{i}$ are assumed to be standardized Gaussian variables with zero mean and unit variance.
The mutual information between $X$ and $Y$ is given by solving the integral in (4) analytically [21]:
$\phantom{.}I(X,Y)=-\frac{d}{2}\ln(1-\rho^{2}).$ (17)
The marginal entropy for a multivariate Gaussian $X$ is [1, Theorem 8.4.1]:
$\phantom{,}H(X)=\frac{1}{2}\ln\left((2\pi\mathrm{e})^{d}|\Sigma_{X}|\right),$ (18)
where $|\Sigma_{X}|$ is the determinant of the covariance matrix $\Sigma_{X}$ and $d$ is the dimensionality of $X$. In the special case where both $X$ and $Y$ are standard normal variables (i.e. $\Sigma_{X}=\Sigma_{Y}=I_{d}$), their entropies simplify to:
$\phantom{.}H(X)=H(Y)=\frac{d}{2}\ln(2\pi\mathrm{e}).$ (19)
By substituting the expressions from (17) and (19) into (6), we obtain the following equation for the true NMI:
$\phantom{.}I_{\text{NMI}}(X,Y)=-\frac{\ln(1-\rho^{2})}{\ln(2\pi\mathrm{e})}.$ (20)
As $\rho$ approaches $1$, $\ln(1-\rho^{2})$ tends to $\infty$- ∞, leading the Equation (20) to diverge. To ensure stability, we cap the NMI at $1.0$, which corresponds to the case of perfect correlation.
In our experiments, we generate 10,000 samples for each case, with $\rho$ values ranging from 00 to $1$ in increments of $0.1$, and dimensionalities $d=1,2,4,8,16,32,64,128,256,512$. The number of neighbors $K$ is set to $5$, a common choice in similar studies. Each experiment is repeated 10 times, with both the means and variances of the results reported.
IV-B Results
Our results (Figure 2) show that the KSG estimator undershoots in low dimensions (1 and 2), overshoots in moderate dimensions (4, 8, 16, and 32), and again undershoots in high dimensions (64, 128, 256, and 512), with the baseline method failing beyond 256 dimensions due to numerical overflow in the scaling-invariant k-NN radius calculation. The proposed transformation resolves this issue, ensuring stability even at 512 dimensions.
In our experiments, dimensionality $d$ has a much greater impact on performance than the correlation coefficient $\rho$. As dimensions increase, the KSG estimator struggles, especially beyond 8 dimensions. While the variance remains generally reasonable, a notable increase occurs at 32 dimensions. Both methods perform similarly overall, but the proposed method resolves numerical issues in high dimensions.
These results align with recent benchmarks showing that KSG performs well in low and moderate dimensions but struggles in high-dimensional and sparse interaction settings [25]. This is likely due to the curse of dimensionality, where data sparsity and complex interactions would require larger sample sizes for accurate estimates.
In summary, while the baseline estimator’s performance is strongly affected by dimensionality, the proposed transformation enhances its stability in high-dimensional scenarios, allowing it to handle cases where it would otherwise fail.
Figure 2: Results of NMI estimation for multivariate Gaussian data with the joint covariance matrix from Equation (16) across varying dimensionalities from $1$ to $512$, plotted against the correlation coefficient $\rho$, with $\rho$ values ranging from 00 to $1$ in increments of $0.1$. Each subplot corresponds to a specific dimensionality, showing true NMI, baseline method estimates, and estimates using the proposed method. The baseline method fails at higher dimensions beyond 256 due to overflow, while the proposed method continues to provide numerically stable results.
V Conclusion
Mutual information is an essential statistical measure for quantifying dependencies in fields like machine learning and statistics. In this work, we addressed a limitation in Normalized Mutual Information (NMI) estimation using the Kraskov-Stögbauer-Grassberger (KSG) estimator, originally proposed by Nagel et al. [20]. Their method suffers from numerical overflow in high-dimensional settings, particularly when calculating the scaling-invariant k-NN radius as joint dimensionality exceeds several hundred, as demonstrated in our experiments.
To resolve this issue, we proposed a logarithmic transformation that mitigates the overflow problem while maintaining precision. This transformation was mathematically proven and empirically validated across various dimensional and correlation settings using multivariate Gaussian data. Our results show that while the baseline method fails in high dimensions, the proposed transformation ensures numerical stability up to 512 dimensions.
Moreover, we found that dimensionality impacted the KSG estimator’s performance more than the correlation coefficient, aligning with recent benchmarks by Czyż et al. [25], and highlighting challenges in high-dimensional, sparse interaction scenarios. We believe further validation of the proposed transformation is unnecessary as the transformation is proven to be correct in the present work and validated experimentally.
In conclusion, the proposed transformation significantly enhances the numerical stability of NMI estimation in high-dimensional settings. It addresses a critical limitation of the baseline method, extending the applicability of KSG-based NMI estimators and providing a more stable tool for practitioners across various fields.